OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 2x − y = 4.

OpenStudy (amistre64):

\[ax+by=c\] \[\frac{ax}{c}+\frac{by}{c}=\frac{c}{c}\] \[\frac{x}{c/a}+\frac{y}{c/b}=1\] when x=0, y = c/b when y=0, x = c/a